The names of 5 men and 5 women are written on slips of paper. Four names are drawn. 

Find the probability using hyper geometric distribution that 2 are men and 2 are 

women?


1
Expert's answer
2021-06-29T04:40:21-0400

Let "X="the number of men names drawn.

"P(X=x)=\dfrac{\dbinom{k}{x}\dbinom{N-k}{n-x}}{\dbinom{N}{n}}"

Given "k=5, N=5+5=10, n=4."


"P(X=2)=\dfrac{\dbinom{5}{2}\dbinom{5}{4-2}}{\dbinom{10}{4}}=\dfrac{10(10)}{210}=\dfrac{10}{21}"